OpenStudy (anonymous):

A building in a city has a rectangular base. The length of the base measures 75ft less than twice the width. The perimeter of this base 810 ft. What are the dimensions of the base?

OpenStudy (anonymous):

given, Length(L)=(2*width(w))-75, ----- (1) Perimeter of rectangle(p)=2(L+w)=810 ----- (2) Sustituting equation 1 in 2 2(2*w-75+w)=810 2(3w-75)=810 3w-75=405 3w=480 w=160 substitute the value of w in in equation#2 2(L+160)=810 L=245 The dimensions of base is 245*160